OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (–1, –4) and (2, 5). Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (amistre64):

slope is important; so lets find it; subtract points and stack y/x (2 , 5) (–1, –4) ------- 3 , 9 ; slope = 9/3 = 3 the equation of the line is: y = mx -mPx + Py , where (Px,Py) is any point that is given

OpenStudy (amistre64):

and m=slope :) y = 3x -3(2)+5 y = 3x -6+5 y = 3x -1

OpenStudy (amistre64):

since we used both points to form the line; it really doesnt matter which one we use to calibrate the equation with